OpenStudy (anonymous):

find the slope and y intercept of the line in the given equation. y=-2/7x-3

OpenStudy (anonymous):

when an eqaution is in this form (y=mx+b), the number multiplying the x is the slope. Y intecept is where x =0. so plug in 0 for x and evaluate

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

so what's the slope?

OpenStudy (anonymous):

would be 7 or 3

OpenStudy (anonymous):

im new to this algebra so Im trying to understand

OpenStudy (australopithecus):

for the slope use this equation slope = m m = (y2 - y1)/(x2-x1) to find the y intercepts set x = 0 and solve for y you will get (0,y) you will only get one or no y-intercepts for the x-intercepts set y = 0 and solve for x you will get (x,0) you can have more than one solution (more than one point) in some cases

OpenStudy (australopithecus):

there solved

OpenStudy (anonymous):

so would I graph this

OpenStudy (australopithecus):

also for the slope equation (x1,y1) and (x2,y2) are two points in the function (equation) you are looking at so for y=-2/7x-3 just sub values into x and solve for y to acquire points then use the equation for the slope I showed you to solve for the slope

OpenStudy (australopithecus):

you dont have to graph it you can just use the formula I provided, sub two numbers into the equation in place of x and solve for y to get two points to find the slope.

OpenStudy (australopithecus):

the equation being what is refereed to as a function

OpenStudy (australopithecus):

referred*

OpenStudy (australopithecus):

Also note in some cases you will have no x intercepts for example 2^(x) = y y can never equal 0 thus you will never have any x intercepts the graph looks like this  it will come infinity close to zero on the x-axis but never reach it

OpenStudy (anonymous):

im sorry im just not sure how to write it